# Quantum accessible information and classical entropy inequalities

### v3

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed. Via the optimality criterion, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
1
Introduction
@@
The proposed entropy inequalities may be more convenient for numerical check than the direct optimization, however proving them analytically is a challenge. In section 5, the suggestions for proofs proposed by the second author as a response to the hypotheses (27), (67), (79) put forward by the first author, are presented. Via the optimality criterion of section 2, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids formulated and numerically substantiated in [2].
2
Accessible information: a criterion for optimality
@@
$A(\mathcal{E})=\sup_{\mathcal{M}}I(\mathcal{E},\mathcal{M}).$
@@
Introducing the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
@@
and observable $\mathcal{M}^{\prime}$ with the components [^1] In the case $\overline{\rho}$ is degenerate, one should use generalized inverse in the following formula.
@@
Ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ with $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|$ is the solution of the problem (5); $\Lambda_{0}$ is the solution of the problem (6);
@@
$\left[K(\sigma)-\Lambda_{0}\right]\,|\phi_{k}^{0}\rangle=0,\,k=1,\dots,n$.
@@
The value of the accessible information (4) is
@@
where $\Lambda_{0}$ must be Hermitean operator.
@@
by necessity form an orthonormal basis. Then the entropy inequality (7) takes the form
@@
where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint $\sum_{j=1}^{d}\left|z_{j}\right|^{2}=1$. The left-hand side is just the Shannon entropy of the p.d. $\left\{\left|z_{j}\right|^{2};j=1,\dots,d\right\}$, and such an inequality can be approached with the tools of classical analysis.
The optimality conditions can be useful in two ways:
@@ 3 Example: two pure states
Multiplying (14) by $\overline{\rho}^{-1/2}$ from the right and taking into account (19) we obtain
$\Lambda_{0}\overline{\rho}^{1/2}$ $=$ $-\frac{1}{\sqrt{2}}\left(\left|e_{+}\rangle\langle e_{+}\right|+\left|e_{-}\rangle\langle e_{-}\right|\right)\sqrt{p}\log p$
$-\frac{1}{\sqrt{2}}\left(\left|e_{+}\rangle\langle e_{-}\right|+\left|e_{-}\rangle\langle e_{+}\right|\right)\sqrt{1-p}\log(1-p)$
$=$ $-\frac{1}{\sqrt{2}}\left|e_{0}\rangle\langle e_{0}\right|\left(\sqrt{p}\log p+\sqrt{1-p}\log(1-p)\right)$
$-\frac{1}{\sqrt{2}}\left|e_{1}\rangle\langle e_{1}\right|\left(\sqrt{p}\log p-\sqrt{1-p}\log(1-p)\right).$
Taking into account (18) and noticing that
$\cos\alpha/2$ $=$ $\frac{1}{\sqrt{2}}\left(\sqrt{p}+\sqrt{1-p}\right)$
$\sin\alpha/2$ $=$ $\frac{1}{\sqrt{2}}\left(\sqrt{p}-\sqrt{1-p}\right),$
we obtain
$\Lambda_{0}=\lambda_{0}(p)\left|e_{0}\rangle\langle e_{0}\right|-\lambda_{1}(p)\left|e_{1}\rangle\langle e_{1}\right|,$ (22)
where
@@
in accordance with (17).
@@ 4.2 The conjecture: “quantum pyramids”
$a=\sqrt{mr_{1}}+\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},\quad b=\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}}.$ (37)
@@
Operators of the form $c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|$ constitute an Abelian algebra.
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 in previous subsection we look for $\Lambda_{0}$ in the form
@@
which is the second line, taking into account that
$1-r_{0}=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}$
by (34), (37). ∎
@@ 4.4 The case of obtuse pyramids
For complex $z_{j}$ satisfying
$\mathit{\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,}$ (66)
it holds
@@
Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$. However, as the following lemma shows, we can always restrict to the case of real $z_{j}$.
Lemma 1.
Let $\tilde{\mu}_{0},\tilde{\mu}_{1}$ be real numbers. If the inequality
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}\leq\tilde{\mu}_{1}$ (70)
If it holds for all real $z_{j}$ satisfying the condition (66) and
$\mathit{\sum_{j=1}^{m}z_{j}=0,}$ (71)
then it holds for all complex $z_{j}$ satisfying the conditions (66), (71).
Proof.
Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (66) amounts to
$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1,\quad$ (72)
and (71) amounts to
$\sum_{j=1}^{m}x_{j}=0,\,\sum_{j=1}^{m}y_{j}=0.$ (73)
Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, hence
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}$ $=$ $\sum_{j=1}^{m}f\left(P\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}+(1-P)\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)$
$+$ $P\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}+(1-P)\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}$
$\leq$ $P\left[\sum_{j=1}^{m}f\left(\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}\right]$
$+$ $(1-P)\left[\sum_{j=1}^{m}f\left(\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}\right]$
$\leq$ $\tilde{\mu}_{1},$
because
and by assumption (70) holds for real $z_{j}$ satisfying (66). If additionally the complex $z_{j}$ satisfy (71) then
$\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}=\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}=0,$
and the second statement of the lemma follows. ∎
@@ 4.5 The case of flat pyramid
for complex $z_{j}$ satisfying (81)  (due to the second statement of lemma 1 one can restrict here to real $z_{j}$ ). Indeed, for $m\leq 6$ we can use the argument leading to inequality (79) with $\sum_{j=1}^{m}z_{j}=0$ and $p=p(m)$. Then (78) implies the first line in the right-hand side of (82). On the other hand, for $m\geq 7$ we should use the inequality (67) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthand side vanish, while $\tilde{\mu}_{1}(p)\$gets the value (74) i.e. the second line in the righthand side of (82), which is greater than 1 for $m\leq 6$, but less for $m\geq 7$. We thus arrive at the following special case of theorem 3:
@@ 5.1 The case of theorem 2
On the boundary of the simplex of probability distributions the function $F(t)$ cannot reach a minimum, since the function $s\mapsto u(t_{*}+sa)$,
@@
for the boundary point $t_{*}$ and the transversal direction $a$ has the derivative -$\infty$ at zero:
@@
where $I_{0}=\{1\leq j\leq m;\ t_{*,j}=0\}$, $C>0$ is some constant. In this case, $a_{i}>0,\ i\in I_{0}$ (direction inside the simplex), so $u^{\prime}(0)=$-$\infty$.
It is necessary to check the inequality (83) at the critical points (see Lemma 2), which leads to the inequality
$\leq$ $-q\log\dfrac{q}{k/m}-(1-q)\log\dfrac{1-q}{1-k/m},$
where $\mu_{0}(p,\alpha)$ and $\mu_{1}(p,\alpha)$ are defined in (84) and (85) below. Lemma 3 states the validity of this inequality. ∎
Consider the functions
for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
The following lemmas lead to the proof of the main theorem.
@@
The critical points $t_{*}=(t_{*,1},...,t_{*,m})$ of the function
$F(t)=H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}$
have the form
$t_{*}=\Bigl{(}q/k,...,q/k,(1-q)/(m-k),...,(1-q)/(m-k)\Bigl{)}$
up to permutations, for some $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$). One of the critical points is equal to (p,(1p)/(m1),,(1p)/(m1))(p,(1-p)/(m-1),...,(1-p)/(m-1)), in which the value of the function is equal to 0.
@@
Instead of $F(t)$ we consider a function of the form $u(s)=F(t+sa)$, where $t=(t_{1},...,t_{m}),\ a=(a_{1},...,a_{m}),\ \sum\limits_{i}a_{i}=0$, on such a range of values of $s$ that $t+as$ defines a probability distribution.
@@
Let us denote $a=\ln t_{*,1}+\dfrac{\mu_{0}(p)\ln(2)}{\sqrt{t_{*,1}}}\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl{)},\ C=\mu_{0}(p)\ln(2)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl{)}$. The equation
$a-\ln(s)=C/\sqrt{s},\quad s>0,$ (86)
has no more than two roots, since the derivatives of the right and left parts with respect to $s$ coincide only when $s=C^{2}/4$. But all the numbers $t_{*,1},...,t_{*,m}$ are solutions of the equation (86). Thus, among the values of $\{t_{*,i}\}$ there are no more than two different values, and the lemma is proved. ∎
The proof of the following lemma is based on the behavior of the difference between binary entropy and a function containing both a linear component and a square root term $\sqrt{t(1-t)}$. We conduct a qualitative analysis of the minimizer of this function. The global minimum (there can be no more than two local minima) appears on either the interval $[0,1/2]$ or $[1/2,1]$. Importantly, this segment contains no other critical points. Therefore, minimizing the function reduces to identifying the critical point within the appropriate interval and evaluating the function at that point.
A detailed proof of this statement will be provided at the end of the section.
The variables $\alpha,\ p,\beta,\ t$ used in what follows, unless otherwise stated, are such that
@@
For introduced $\alpha,\ p,\ \beta,\ t$ the inequality
@@
$\leq$ $-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}$
holds.
@@
$f(t)=-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigl{(}\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)\Bigl{)},$
for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (89) (see lemma 4 below).
The derivative of $f(t)$ has form
@@
    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to the lemma 4 presented later. One has
@@
    Therefore, $V(p)\geq 0$ and the inequality (88) is satisfied.
@@
    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$. Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$. We obtained the desired inequality $f(t)\geq 0$ for the minimizer $t_{*}$ and it is even more so for the other points $t\in[0,1]$.
In the particular case of (88), where $\alpha=1/2$, one has an interesting inequality
@@
The proof of lemma 3 is based on a simple fact of analysis:
Lemma 4.
Let $A>0,\ B,\ C\in\mathbb{R}$ (The notation of the coefficients is used only within the framework of the formulation of the lemma and its proof). Then the function
$g(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (89)
[…]
Proof.
$g^{\prime}(0)=-\infty,\quad g^{\prime}(1)=+\infty,$
therefore local minima are on the interval $(0,1)$.
In the calculations we will move on to the natural logarithm.
$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
$g^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
$g^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
Therefore, there are two inflection points located on different sides of $t=1/2$. It follows that there are 3 critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.

### v4

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. We show that the recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed. Via the optimality criterion, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
1 Introduction
@@
The proposed entropy inequalities may be more convenient for numerical check than the direct optimization, however proving them analytically is a challenge. In section 5, the suggestions for proofs proposed by the second author as a response to the hypotheses (31), (72), (80) put forward by the first author, are presented. Via the optimality criterion of section 2, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids formulated and numerically substantiated in [2].
2 Accessible information: a criterion for optimality
@@
$A(\mathcal{E})=\sup_{\mathcal{M}}I(\mathcal{E},\mathcal{M}).$ (1)
@@
From now on we assume that the average state $\overline{\rho}$ is nondegenerate. Introducing the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
@@
and observable $\mathcal{M}^{\prime}$ with the components
@@
$\Lambda_{0}$ is the solution of the problem (7); ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ with $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|$ is the solution of the problem (6) (equivalently, observable $M_{k}^{0}=\left|\varphi_{k}^{o}\rangle\langle\varphi_{k}^{0}\right|,\,k=1,\dots,n,$ with
$|\varphi_{k}^{0}\rangle=\left(p_{k}^{0}\right)^{1/2}\overline{\rho}^{-1/2}|\phi_{k}^{0}\rangle,$
maximizes the accessible information (1));
@@
$\left[K(\sigma_{k}^{0})-\Lambda_{0}\right]\,|\phi_{k}^{0}\rangle=0,\,k=1,\dots,n$.
@@
The solution of the dual problem (7) is unique on the support of the operator $\overline{\rho}$ i.e. $\Lambda_{0}\overline{\rho}=\Lambda_{0}^{\prime}\overline{\rho}$ for any two solutions $\Lambda_{0},\Lambda_{0}^{\prime}.$ In particular, if $\overline{\rho}$ is non-degenerate, then the solution of the dual problem is unique.
According to (5), the value of the accessible information (5) is then
@@
where $\Lambda_{0}$ must be Hermitean operator. This implies that $\Lambda_{0}\overline{\rho}$ is the same for all solutions of the dual problem.
@@
by necessity form an orthonormal basis (depending on the ensemble $\mathcal{E}$). Then the entropy inequality (8) takes the form
@@
where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1.$ (18)
The left-hand side is just the Shannon entropy of the p.d. $\left\{\left|z_{j}\right|^{2};j=1,\dots,d\right\}$, and such an inequality can be approached with the tools of classical analysis. We will need the following
Lemma 1.
Let $\lambda_{jk}$ be real numbers. If the inequality
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k}\leq 0$ (19)
Proof.
Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (18) amounts to
$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1.\quad$ (20)
Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, and $\lambda_{jk}$ are real hence
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k}$ $=$ $\sum_{j=1}^{m}f\left(P\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}+(1-P)\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)$
$+$ $P\sum_{j=1}^{d}\lambda_{jk}x_{j}x_{k}+(1-P)\sum_{j=1}^{d}\lambda_{jk}y_{j}y_{k}$
$\leq$ $P\left[\sum_{j=1}^{m}f\left(\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}\right)+\sum_{j=1}^{d}\lambda_{jk}x_{j}x_{k}\right]$
$+$ $(1-P)\left[\sum_{j=1}^{m}f\left(\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)+\sum_{j=1}^{d}\lambda_{jk}y_{j}y_{k}\right]\leq 0,$
because
and by assumption (19) holds for real $z_{j}$ satisfying (18). ∎
The optimality conditions can be useful in the following ways:
@@
  3. 3.
    The uniqueness of the solution of the dual problem (assuming $\bar{\rho}$ non-degenerate) has the following useful consequence. Assume that the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ is symmetric in that $\pi_{j}\equiv 1/m$ and there is a group of *-automorphisms $\left\{\alpha_{g};g\in G\right\}$ of the algebra of operators on $\mathcal{H}$ acting transitively on the set of states $\left\{\rho_{j}\right\}_{j=1,\dots,m}$. Then $\Lambda_{0}$ is an invariant of this group: $\alpha_{g}^{\ast}\left[\Lambda_{0}\right]=\Lambda_{0};g\in G.$
@@ 3 Example: two pure states
The ensemble $\mathcal{E}$ is symmetric with respect to the reflection relative to the axis $|e_{0}\rangle;$ according to the remark 3 at the end of previous section, $\Lambda_{0}$ should also obey this symmetry, hence
$\Lambda_{0}=\lambda_{0}(p)\left|e_{0}\rangle\langle e_{0}\right|-\lambda_{1}(p)\left|e_{1}\rangle\langle e_{1}\right|.$ (26)
Solving the conditions (ii.b)
$\left[K(\sigma_{\pm}^{0})-\Lambda_{0}\right]\,|\phi_{\pm}^{0}\rangle=0$
and taking into account that
$\cos\alpha/2=\frac{1}{\sqrt{2}}\left(\sqrt{p}+\sqrt{1-p}\right),\quad\sin\alpha/2=\frac{1}{\sqrt{2}}\left(\sqrt{p}-\sqrt{1-p}\right),$
results in
@@
and (10) implies $A(\mathcal{E})=1-h(p)$ in accordance with (21).
@@ 4.2 The conjecture: “quantum pyramids”
$a=\sqrt{mr_{1}}+\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},\quad b=\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},$ (41)
$1-r_{0}=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}.$ (42)
@@
Operators of the form $c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|$ constitute Abelian algebra of operators invariant under the group of all *-automorphisms generated by the orthogonal transformations leaving the quantum pyramid invariant.
@@
Remark 3.
An upper bound and a benchmark for $A(\mathcal{E})$ is the classical capacity of the c-q channel $j\rightarrow|\psi_{j}\rangle\langle\psi_{j}|$:
$C=\max_{\pi}H\left(\sum_{j=1}^{m}\pi_{j}|\psi_{j}\rangle\langle\psi_{j}|\right)=H(\bar{\rho}).$ (51)
The last equality follows from the symmetry of the quantum pyramid and concavity of the von Neumann entropy. According to (43) the density operator $\bar{\rho}$ has the eigenvalues $r_{0},r_{1}$, the last of multiplicity m1m-1, hence
$C$ $=$ $-r_{0}\log r_{0}-(m-1)r_{1}\log r_{1}$
$=$ $-\left(1-\frac{m-1}{m}(a-b)^{2}\right)\log\left(1-\frac{m-1}{m}(a-b)^{2}\right)$
- $\frac{m-1}{m}(a-b)^{2}\log\frac{(a-b)^{2}}{m},$
where in the second equality we used (42).
Unlike the case of accessible information, this expression is the same for all values of the parameters, for acute, obtuse and flat pyramids. The ratio $C/A(\mathcal{E})$ gives the gain due to the measurement entanglement at the channel’s output. In particular, it tends to the infinity as the pyramid collapses to ort (weak signal) and equals to 1 for rectangular pyramid (ideal channel).
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 we look for $\Lambda_{0}$ in the form
@@
which is the second line, taking into account (42). ∎
@@ 4.4 The case of obtuse pyramids
For complex $z_{j}$ satisfying $\mathit{\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,}$ it holds
@@
Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$. However, as the lemma 1 shows, we can always restrict to the case of real $z_{j}$.
@@ 4.5 The case of flat pyramid
for complex $z_{j}$ satisfying (82) (due to a simple modification of lemma 1 one can restrict here to real $z_{j}$ ). Indeed, for $m\leq 6$ we can use the argument leading to inequality (80) with $\sum_{j=1}^{m}z_{j}=0$ and $p=p(m)$. Then (79) implies the first line in the right-hand side of (83). On the other hand, for $m\geq 7$ we should use the inequality (72) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthand side vanish, while $\tilde{\mu}_{1}(p)\$gets the value (75) i.e. the second line in the righthand side of (83), which is greater than 1 for $m\leq 6$, but less for $m\geq 7$. We thus arrive at the following special case of theorem 3:
@@ 5.1 The case of theorem 2
The function $F(t)$ cannot reach a minimum on the boundary of the simplex of probability distributions, since the function
@@
for a boundary point $t_{*}$ and a transversal direction $a$ has the derivative -$\infty$ at zero:
@@
where $I_{0}=\{1\leq j\leq m;\ t_{*,j}=0\}$, $C>0$ is a constant. In this case, $a_{i}>0,\ i\in I_{0}$ (direction inside the simplex), hence $u^{\prime}(0)=$-$\infty$.∎
Thus it is sufficient to check the inequality (84) only at the critical points.
@@
The critical points $t_{*}=(t_{*,1},...,t_{*,m})$ of the function $F(t)$ have the form
$t_{*}=\Bigl(q/k,...,q/k,(1-q)/(m-k),...,(1-q)/(m-k)\Bigl)$ (85)
up to permutations, for some integer $k$, $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$). One of the critical points is (p,(1p)/(m1),,(1p)/(m1))(p,(1-p)/(m-1),...,(1-p)/(m-1)), at which the value of the function $F(t)$ is equal to 0.
@@
Consider the function $u(s)=F(t+sa)$, where $t=(t_{1},...,t_{m}),\ a=(a_{1},...,a_{m}),\ \sum\limits_{i}a_{i}=0$, for a range of $s$ such that $t+as$ defines a probability distribution. Then
@@
Let us denote $C_{1}=\mu_{0}(p)\Bigl(\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl),C_{2}=\log t_{*,1}+\dfrac{C_{1}}{\sqrt{t_{*,1}}}$. The equation
$C_{2}-\log s=C_{1}/\sqrt{s},\quad s>0,$ (86)
has at most two roots, since the derivatives of the right and the left parts with respect to $s$ coincide only when $s=C_{1}^{2}/4$. But all the numbers $t_{*,1},...,t_{*,m}$ are solutions of the equation (86). Thus, among the values of $\{t_{*,i}\}$ there are no more than two different values, hence the lemma follows. ∎
At the critical points (85) the inequality $F(t_{*})\geq 0$ reduces to
$\leq$ $-q\log\dfrac{q}{k/m}-(1-q)\log\dfrac{1-q}{1-k/m},$
where $\mu_{0}(p,\alpha)$ and $\mu_{1}(p,\alpha)$ are defined as
for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
But this inequality is a special case of lemma 4 below, the proof of which is based on the behavior of the difference between binary entropy and a function containing both a linear component and a square root term $\sqrt{t(1-t)}$. We perform a qualitative analysis of the minimizer of this function. The global minimum (there can be no more than two local minima) appears on either the interval $[0,1/2]$ or $[1/2,1]$. Importantly, this segment contains no other critical points. Therefore, minimizing the function reduces to identifying the critical point within the appropriate interval and evaluating the function at that point. A detailed proof of this statement follows from
Lemma 3.
Let $A>0,\ B,\ C\in\mathbb{R}$ (the notation of the coefficients is used only within the framework of this lemma). Then the function
$g(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (89)
[…]
Proof.
We have
$g^{\prime}(0)=-\infty,\quad g^{\prime}(1)=+\infty,$
therefore local minima are on the interval $(0,1)$.
In the calculations we will pass to the natural logarithm.
$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
$g^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
$g^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
Therefore, there are two inflection points located on different sides of $t=1/2$. It follows that there are three critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.
Let us finally proceed to the proof of lemma 4. The variables $\alpha,\ p,\ \beta,\ t$ used in what follows, unless otherwise stated, are such that
@@
For $\alpha,\ p,\ \beta,\ t$ specified above one has the inequality
@@
$\leq$ $-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}.$
@@
$f(t)=-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigl[\mu_{0}(p,\alpha)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\mu_{1}(p,\alpha)\Bigl],$
for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (89) of lemma 3.
The derivative of $f(t)$ is
@@
    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to lemma 3 presented later. One has
@@
    Therefore, $V(p)\geq 0$ and the inequality (91) is established.
@@
    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$. Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$. We obtained the desired inequality $f(t)\geq 0$ for the minimizer $t_{*}$ and hence for all the other points $t\in[0,1]$.
In the special case of (91), where $\alpha=1/2$, one has an interesting inequality